OpenStudy (anonymous):

y=-x^2-6x interval when function is increasing? interval when function is decreasing?

OpenStudy (anonymous):

The derivative is \[y \prime = -2x-6\] When \[y \prime > 0\] The function is increasing If \[y \prime < 0\] The function is decreasing, then?...

OpenStudy (anonymous):

The derivative changes in x=-3 So is increasing in the interval\[(-\infty, -3)\] (Because y'>0 or positive) and decreasing in \[(-3,\infty)\] (because y' is negative)

OpenStudy (anonymous):

Thank you.. how to do I fiqure out the vertex

OpenStudy (anonymous):

Complete the polynomial, to find the vertex

OpenStudy (anonymous):

\[y=-x ^{2}-6x\] \[y=-(x ^{2}+6x)\] Rounding out the perfect square trinomial (I used google traductor) \[y=-\left[ \left( x+3 \right)^{2}-9 \right]\] \[y=-(x+3)^{2}+9\] \[y-9=-(x+3)^{2}\] From this equation we can gen the vertex V(-3,9)